### v1

# Under the null of valid specification, pre-tests of valid specification do not distort inference

Abstract
Consider a parameter of interest, which can be consistently estimated under some conditions. Suppose also that we can at least partly test these conditions with specification tests. We consider the common practice of conducting inference on the parameter of interest conditional on not rejecting these tests. We show that if the tested conditions hold, conditional inference is valid but possibly conservative. This holds generally, without imposing any assumption on the asymptotic dependence between the estimator of the parameter of interest and the specification test.
1 Introduction
Theoretical guarantees on inference methods are usually given unconditionally. Yet, inference on a parameter of interest is rarely conducted in isolation of other statistical practices. In this paper, we consider the common practice of reporting confidence intervals or performing tests on a parameter of interest, $\beta_{0}$, only if specification tests are not rejected. For instance, one may proceed with a difference-in-difference (DID) estimator only if a pretrends test is not rejected. One may then wonder if pre-testing may distort inference. Let $\widehat{\beta}$ denote an estimator of $\beta_{0}$. This note shows that under the null that the conditions underlying the validity of $\widehat{\beta}$ hold, conditional inference is valid, albeit possibly conservative. This result holds under general conditions. In particular, no restriction on the asymptotic dependence between $\widehat{\beta}$ and the specification tests is needed. The main condition on these specification tests is that they are convex functionals. This holds for finite-dimensional $F$-tests but also for infinite-dimensional tests based on empirical processes such as (weighted) Kolmogorov-Smirnov or Crámer-von Mises tests. For finite-dimensional test statistics, our result is a direct application of the Gaussian correlation inequality (Royen, 2014). For infinite-dimensional ones, we first rely on an approximation before applying this inequality.
For instance, in a DID analysis, our result implies that under the null of parallel trends, inference conditional on not rejecting a pretrends test is at worst conservative, but can never be liberal, irrespective of the correlation structure between the DID and pre-trends estimators. This has implications for some recent discussions in that literature. For instance, with staggered treatment timing, valid inference conditional on not rejecting pre-trends tests under the null of no pre-trends was put forward as an argument in favor of some of the recently proposed heterogeneity-robust estimators (Borusyak et al., 2024). Actually, our result implies that all heterogeneity-robust estimators lead to exact or conservative inference conditional on not rejecting a pretrends test.
Importantly, our results only hold under the null that the conditions underlying the validity of $\widehat{\beta}$ hold. Thus, they have no bearing on the negative results of Leeb and Pötscher (2005) and Leeb and Pötscher (2006), who show that under local violations of those conditions, pre-testing can distort inference. Under the alternative that the assumptions underlying it are violated, $\widehat{\beta}$ is biased and leads to invalid inference anyways, so the question becomes whether pre-testing, and using $\widehat{\beta}$ when the pre-test is incorrectly not rejected, leads to a larger bias and size distortion than not pre-testing and using $\widehat{\beta}$ anyways. There is probably not a universal answer to that question. In the DID example, Roth (2022) conducts extensive simulations taylored to a large number of published articles in economics. His results suggest on average a small bias aggravation due to pre-testing, though there are also examples where pre-testing reduces the estimator’s bias.
2 The set-up and main result
We are interested in inference on a parameter $\beta_{0}\in\mathbb{R}^{p}$. To this end, we rely on an estimator $\widehat{\beta}$ which is supposed to be consistent for $\beta_{0}$, but only if some conditions are satisfied. We assume to have at our disposal one or several statistical tests of either these conditions themselves, or implications of these conditions. The question we ask is whether inference on $\beta_{0}$ based on $\widehat{\beta}$ is valid when it is performed only conditional on not rejecting any of these tests.
Example 1 (Difference-in-differences)
We assume that some units (the “treated”) receive a binary treatment at a period of time, whereas other units (“the controls”) do not. We are interested in $\beta_{0}$, the average treatment effect on the treated after treatment occurs. $\beta_{0}$ could also correspond to the vector of average treatment effects at different horizons after the treatment takes place. If parallel trends hold, $\beta_{0}$ can be consistently estimated by $\widehat{\beta}$, the usual difference-in-difference estimator. Parallel trends at all periods also imply that $\theta_{0}$, a difference-in-differences estimand on pre-treatment periods, should be equal to 0. We report $\widehat{\beta}$ and confidence intervals on $\beta_{0}$ conditional on not rejecting $\theta_{0}=0$. Note that a common practice is to test separately $\theta_{01}=0$, $\theta_{02}=0$ until $\theta_{0q}=0$. Then, we make inference on $\beta_{0}$ conditional on not rejecting any of these tests.
Example 2 (Instrumental variables)
We seek to estimate $\beta_{0}$, the local average treatment effect of an endogenous binary treatment $D$, using a binary instrumental variable $Z$. Under instrument independence, exclusion restriction and instrument monotonicity, $\beta_{0}$ is equal to the Wald estimand (Imbens and Angrist, 1994). We can test these conditions using for instance Kitagawa’s variance-weighted Kolmogorov-Smirnov test statistic (Kitagawa, 2015). Oftentimes, researchers also perform balancing tests on covariates, as a proxy for testing the instrument independence.
Example 3 (GMM)
We are interested in $\beta_{0}\in\mathbb{R}^{p}$, which satisfies the moment conditions $E[g(U,\beta_{0})]=0$, where $U$ denotes an observed random vector, and $g(u,\beta)\in\mathbb{R}^{q}$, $q\geq p$. If $q>p$, we can test these conditions with an overidentifying test, using either the standard $J$-test or a weighting matrix different from the asymptotically optimal one. In the latter case, the test statistic does not converge to a chi-squared distribution, but we can still perform the test using the bootstrap, for instance. We perform inference in $\beta_{0}$ only when we do not reject the overidentifying test.
Example 4 (Linear and constant effect models with uncounfoundedness)
Consider a continuous treatment $D$ with support $\mathcal{D}$, let $Y(d)$ denote the potential outcome corresponding to $D=d$ and let $Y:=Y(D)$ be the observed outcome. Assume that i) $E(Y(0)|D)=\alpha_{0}$ and ii) $d\mapsto Y(d)=Y(0)+\beta_{0}d$, for two real numbers $(\alpha_{0},\beta_{0})$. Then, $\beta_{0}$ can be estimated by a simple linear regression of $Y$ on $D$, and let $\widehat{\beta}$ denote the corresponding OLS estimator. i) and ii) imply that $E[Y|D]=\alpha_{0}+\beta_{0}D$, a testable implication (de Chaisemartin and D’Haultfœuille, 2024). Define the function $\theta_{0}$ by
$\theta_{0}(d)=E[\mathds{1}\left\{D\leq d\right\}(Y-\alpha_{0}-\beta_{0}D)],$
where $\alpha_{0}=E[Y]-\beta_{0}E[D]$. If $D\perp\!\!\!\perp(Y(d))_{d\in\mathcal{D}}$ and $d\mapsto Y(d)$ is linear, then $d\mapsto E[Y|D=d]$ is linear, which is equivalent to having $\theta_{0}(d)=0$ for all $d\in\mathcal{D}$.
We formalize our problem with the following assumption. Hereafter, we allow for test statistics that are functionals of the empirical measure. We thus use the corresponding standard terminology and notation below; see, e.g., Van der Vaart and Wellner (2023).
Assumption 1
  1. 1.
    We observe a sample $(U_{i})_{i=1,...,n}$ of i.i.d. random vectors with probability distribution $P_{U}$.
  2. 2.
    If $P_{U}\in\mathcal{P}_{0}$, we have
    $\sqrt{n}\left(\widehat{\beta}-\beta_{0}\right)=n^{-1/2}\sum_{i=1}^{n}\psi(U_{i})+o_{P}(1),$ (1)
    where $E[\psi(U_{1})]=0$ and $E[\|\psi(U_{1})\|^{2}]<\infty$ and $V(\psi(U_{1}))$ is nonsingular.
  3. 3.
    If $P_{U}\in\mathcal{P}_{1}$, with $\mathcal{P}_{0}\subseteq\mathcal{P}_{1}$, we have $Pf:=E[f(U_{1})]=0$ for all $f\in\mathcal{F}$, a Donsker class of real-valued functions. [^1] For the definition of a Donsker class, see, e.g., Van der Vaart and Wellner (2023), p.130.
  4. 4.
    We consider $J$ tests of $H_{0}:\;P_{U}\in\mathcal{P}_{1}$ based on the statistics $(T_{1,n},...,T_{J,n})\in\mathbb{R}^{J}$ that satisfy, under $H_{0}$,
    $T_{j,n}=T_{j}(n^{1/2}P_{n})+o_{P}(1),$ (2)
    where $P_{n}$ denotes the empirical measure of $U$, indexed by $\mathcal{F}$. Moreover, $T_{j}$ is convex and satisfies $T_{j}(-P)=T_{j}(P)\geq 0$ and $T_{j}(0)=0$. Finally, $T_{j}$ is continuous with respect to the uniform norm $\|P\|_{\mathcal{F}}:=\sup_{f\in\mathcal{F}}|Pf|$.
  5. 5.
    The critical region of the $j$-th test of $H_{0}$ is $\{T_{j,n}>q_{j,n}\}$, where $q_{j,n}$ is a random variable satisfying $q_{j,n}\stackrel{{\scriptstyle p}}{{\longrightarrow}}q_{j}>0$.
Condition 1 is assumed for simplicity: our results would go through as long as we can establish a uniform central limit theorem. Equation (1) holds under mild regularity restrictions for large classes of estimators (e.g., M- or GMM estimators). Condition 3 states that we can test $H_{0}:\;P_{U}\in\mathcal{P}_{1}$ by checking that some moment conditions hold. Importantly, the class $\mathcal{F}$ may be infinite (and uncountable), in which case we seek to test infinitely many moment restrictions. Conditions 4 and 5 impose some restrictions on the type of tests we are considering. Continuity is typically imposed or checked to derive the weak convergence of $T(P_{n})$ using the continuous mapping theorem. Finally, the last condition imposes the standard condition that the thresholds used in the $J$ tests converge to positive values. The condition $q_{j}>0$ simply rules out that the threshold is at the boundary of the support of the asymptotic distribution of $T_{j,n}$ under $H_{0}$.
We give below three common examples of tests satisfying Conditions 4-5.
Example 5
(F-tests) If $\mathcal{F}$ is finite, =$\mathcal{F}=\{f_{1},...,f_{K}\}$ is Donsker if and only if the functions $f_{1},...,f_{K}$ are measurable and $\max_{k=1,...,K}E[f_{k}(U_{1})^{2}]<\infty$. Let $f=(f_{1},...,f_{K})^{\prime}$ and
$T_{1,n}=n\left(\frac{1}{n}\sum_{i=1}^{n}f(U_{i})\right)^{\prime}\left(\frac{1}{n}\sum_{i=1}^{n}f(U_{i})f(U_{i})^{\prime}\right)^{-1}\left(\frac{1}{n}\sum_{i=1}^{n}f(U_{i})\right).$
We reject the null hypothesis if $T_{1,n}>q_{1-\alpha}(K)$, the quantile of order $1-\alpha$ of a chi-squared distribution with $K$ degrees of freedom. By Slutsky’s lemma,
$T_{1,n}=\left(n^{1/2}(P_{n}f_{1},...,P_{n}f_{K})\right)E[f(U_{1})f(U_{1})^{\prime}]^{-1}\left(n^{1/2}(P_{n}f_{1},...,P_{n}f_{K})^{\prime}\right)+o_{P}(1).$
Therefore, $T_{1,n}=T_{1}(n^{1/2}P_{n})+o_{P}(1)$, where $T_{1}(Q)=(Qf_{1},...,Qf_{K})E[f(U_{1})f(U_{1})^{\prime}]^{-1}(Qf_{1},...,Qf_{K})^{\prime}$. It is easily seen that $T_{1}$ satisfies Condition 4 of Assumption 1.
Example 6
(Weighted Kolmogorov-Smirnov test statistics) Such test statistics are defined by $T_{1}(P)=\sup_{f\in\mathcal{F}}w(f)|Pf|$ for some Donsker clas $\mathcal{F}$. This includes for instance the test statistic of Kitagawa (2015), in which $\mathcal{F}$ is the set of intervals of the real line. [^2] In his case, $T_{1,n}=\sup_{f\in\mathcal{F}}w_{n}(f)|Pf|$ for a random $w_{n}(\cdot)$, but $w_{n}$ converges uniformly to some deterministic $w$, so (3) holds in his context. If we assume that for all $f\in\mathcal{F}$, $0<\underline{w}\leq w(f)\leq\overline{w}<\infty$, then Condition 4 of Assumption 1 holds. To see that it is continuous, fix $P$ and $\delta>0$ and let $Q$ be such that $\|P-Q\|_{\mathcal{F}}<\delta$. Let $f^{*}$ be such that $w(f^{*})|Pf^{*}|>T_{1}(P)-\delta$. Then
$T_{1}(P)-T_{1}(Q)$ $<w(f^{*})|Pf^{*}|-w(f^{*})|Qf^{*}|+\delta$
$\leq\overline{w}\;\sup_{f\in\mathcal{F}}|(P-Q)(f)|+\delta$
$<(\overline{w}+1)\;\delta.$
Example 7
[…]
$|(Pf_{x})^{2}-(QPf_{x})^{2}|$ $=\left|(Pf_{x}+Qf_{x})\times(P-Q)f_{x}\right|$
$\leq\left(\|P\|_{\mathcal{F}}+\|Q\|_{\mathcal{F}}\right)\|P-Q\|_{\mathcal{F}}$
$<\left(2\|P\|_{\mathcal{F}}+\delta\right)\delta.$
Then,
[…]
Theorem 1
Suppose that Assumption 1 holds and $P_{U}\in\mathcal{P}_{0}$. Then, for any convex set $C$ that is symmetric around the origin,
$\lim_{n}P\left[\sqrt{n}\big{(}\widehat{\beta}-\beta_{0}\big{)}\in C|T_{1,n}\leq q_{1,n},...,T_{J,n}\leq q_{J,n}\right]\geq P(Z\in C),$
where $Z\sim\mathcal{N}(0,V(\psi))$. [^3] Assumption 1 does not guarantee that the test statistics $T_{j,n}$ are measurable. If not, the result holds replacing $P$ by the outer probability $P^{*}$ (see, e.g. Van der Vaart and Wellner, 2023, p.6).
The result relies on two main ingredients. To better understand them, suppose for simplicity that $J=1$. The first main ingredient is the approximation of the limit distribution of $T_{1}(n^{1/2}P_{n})$ by a convex function of a finite-dimensional Gaussian measure. The second ingredient is the Gaussian correlation inequality (Royen, 2014), which states that for any centered Gaussian distribution $\mu$ on $\mathbb{R}^{q}$ and two convex sets $E$ and $F$ that are symmmetric about the origin,
$\mu(E\cap F)\geq\mu(E)\mu(F).$
The main implication of Theorem 3 concerns (conditional) inference on $\beta_{0}$. Let $\widehat{V}$ denote an estimator $V(\psi(U_{1}))$. We consider the usual F-test of $\beta_{0}=b_{0}$ with test statistic $F_{n}(b_{0}):=(\widehat{\beta}-b_{0})^{\prime}\widehat{V}^{-1}(\widehat{\beta}-b_{0})$ and critical region $\{F_{n}(b_{0})>q_{1-\alpha}(p)\}$, with $q_{1-\alpha}(p)$ the quantile of order $1-\alpha$ of a chi-squared distribution with $p$ degrees of freedom. We also consider the standard confidence region
$\text{CR}_{1-\alpha}=\left\{b\in\mathbb{R}^{p}:n(\widehat{\beta}-b)^{\prime}\widehat{V}^{-1}(\widehat{\beta}-b)\leq q_{1-\alpha}(p)\right\}.$
Corollary 1
Suppose that Assumption 1 holds, $P_{U}\in\mathcal{P}_{0}$ and $\widehat{V}\stackrel{{\scriptstyle p}}{{\longrightarrow}}V(\psi(U_{1}))$. Then:
  1. 1.
    If $\beta_{0}=b_{0}$, $\lim_{n\to\infty}P\left(F_{n}(b_{0})>q_{1-\alpha}(p)|T_{1,n}\leq q_{1,n},...,T_{J,n}\leq q_{J,n}\right)\leq\alpha$.
  2. 2.
    $\lim_{n\to\infty}P\left(\beta_{0}\in\text{CR}_{1-\alpha}|T_{1,n}\leq q_{1,n},...,T_{J,n}\leq q_{J,n}\right)\geq 1-\alpha$.
Hence, under Assumption 1 and provided that $P_{U}\in\mathcal{P}_{0}$, conditional on accepting specification tests the usual $F$-tests and confidence regions are asymptotically conservative.
3 Extensions
While quite general, the previous set-up excludes some important cases, that we cover now under the following alternate assumption.
Assumption 2
  1. 1.
    We observe a sample $(U_{i})_{i=1,...,n}$ of identically distributed random vectors with probability distribution $P_{U}$.
  2. 2.
    If $P_{U}\in\mathcal{P}_{0}$, we have,
    $\left(\widehat{\Sigma}_{\beta}^{-1/2}\left(\widehat{\beta}-\beta_{0}\right),\widehat{\Sigma}_{\theta}^{-1/2}\left(\widehat{\theta}-\theta_{0}\right)\right)\stackrel{{\scriptstyle d}}{{\longrightarrow}}\mathcal{N}\left(0,\begin{pmatrix}I_{p}&\Sigma_{12}\\ \Sigma_{12}&I_{q}\end{pmatrix}\right),$
    where $I_{k}$ is the identity matrix of size $k$.
  3. 3.
    If $P_{U}\in\mathcal{P}_{1}$, with $\mathcal{P}_{0}\subseteq\mathcal{P}_{1}$, we have $\theta_{0}=0$.
  4. 4.
    We consider $J$ tests of $H_{0}:\;P_{U}\in\mathcal{P}_{1}$ based on the statistics $(T_{1,n},...,T_{J,n})\in\mathbb{R}^{J}$ that satisfy, under $H_{0}$,
    $T_{j,n}=T_{j}\left(\widehat{\Sigma}_{\theta}^{-1/2}\left(\widehat{\theta}-\theta_{0}\right)\right)+o_{P}(1),$ (3)
    where $P_{n}$ denotes the empirical measure of $U$, indexed by $\mathcal{F}$. Moreover, $T_{j}$ is convex on $\mathbb{R}^{q}$ and satisfies $T_{j}(-x)=T_{j}(x)\geq 0$ for all $x\in\mathbb{R}^{q}$ and $T_{j}(0)=0$.
  5. 5.
    The critical region of the $j$-th test of $H_{0}$ is $\{T_{j,n}>q_{j,n}\}$, where $q_{j,n}$ is a random variable satisfying $q_{j,n}\stackrel{{\scriptstyle p}}{{\longrightarrow}}q_{j}>0$.
Unlike Assumption 1, Assumption 2 allows for estimators that converge at rates lower than $n^{1/2}$. An important example is regression discontinuity designs, where specification tests would either be tests that the density is continuous at the threshold (McCrary, 2008) or tests that predetermined variables have continuous means at the threshold. Assumption 2 allows for this case, when the testable implications are of finite dimension, denoted by $q$. Assumption 2 also allows for non-independent but identically distributed observations, as is for instance the case with stationary time series, provided we can prove the asymptotic normality of the actual and specification-test estimators.
Theorem 2
Suppose that Assumption 2 holds and $P_{U}\in\mathcal{P}_{0}$. Then, for any convex set $C$ that is symmetric around the origin,
$\lim_{n}P\left[\widehat{\Sigma}_{\beta}^{-1/2}\left(\widehat{\beta}-\beta_{0}\right)\in C|T_{1,n}\leq q_{1,n},...,T_{J,n}\leq q_{J,n}\right]\geq P(Z\in C),$
where $Z\sim\mathcal{N}(0,I_{p})$.
As above, this theorem implies conservative tests and confidence regions on $\beta_{0}$ conditional on not rejecting specification tests.
Appendix A Proofs
A.1 Proof of Theorem 3
Let us define $Z_{n}:=\sqrt{n}(\widehat{\beta}-\beta_{0})$, $T_{n}:=\max_{j=1,...,J}[T_{j,n}-q_{j,n}]$ and $T(P):=\max_{j=1,...,J}[T_{j}(P)-q_{j}]$. $T$ is convex and continuous, as the maximum of $J$ convex and continuous functionals. Moreover,
$P(Z_{n}\in C,T_{1,n}\leq q_{1,n},...,T_{J,n}\leq q_{J,n})=P(Z_{n}\in C,T_{n}\leq 0).$
By Assumptions 1.3, 1.4 and 1.5
$(Z_{n},T_{n})=(G_{n}\psi,T(n^{1/2}(P_{n}-P))+o_{P}(1).$ (4)
Let $\mathcal{G}=\mathcal{F}\cup\{\psi^{1},...,\psi^{d}\}$. As union of Donsker classes, $\mathcal{G}$ is Donsker (see Van der Vaart and Wellner, 2023, p.136). Hence, the empirical process $G_{n}:=n^{1/2}(P_{n}-P)$ indexed by $\mathcal{G}$ is asymptotically Gaussian; we denote by $G$ its limit. By (4) and the continuous mapping theorem,
$(Z_{n},T_{n})\stackrel{{\scriptstyle d}}{{\longrightarrow}}(G\psi,T(G)).$
Let $\partial A$ denote the boundary of the set $A$. The boundary of $C\times(-\infty,0]$ satisfies
$\partial(C\times(-\infty,0])=(\partial C\times(-\infty,0])\cup(C\times\{0\}).$
Moreover, since $C$ is convex, $\partial C$ has Lebesgue measure zero (see, e.g. Lang, 1986, p.90). Since $V(\psi(U_{1}))$ is nonsingular, the distribution of $G\psi$ is absolutely continuous with respect to the Lebesgue measure. Hence, $P(G\psi\in\partial C)=0$. By Theorem 11.1 and Problem 11.3 in Davydov et al. (1998) and continuity of $T$, the cumulative distribution function $H$ of $T(G)$ is strictly increasing on $(r_{0},\infty)$, with $r_{0}:=\inf_{P}T(P)=\max_{j=1,...,J}-q_{j}<0$. Hence, $H$ is continuous at 0, $H(0)>0$ and $P(T(G)=q)=0$. Hence,
$P((G\psi,T(G))\in\partial(C\times(-\infty,0]))=0.$
Thus, by Portmanteau’s lemma (see, e.g., Lemma 18.9 in van der Vaart, 2000),
$\lim_{n}P^{*}(Z_{n}\in C,T_{n}\leq 0)=P(G\psi\in C,T(G)\leq 0),$
where $P^{*}$ denotes the outer probability. Similarly, $\lim_{n}P^{*}(T_{n}\leq 0)=P(T(G)\leq 0)=H(0)>0$. Hence,
$\lim_{n}P^{*}(Z_{n}\in C|T_{n}\leq 0)=P(G\psi\in C|T(G)\leq 0).$ (5)
Now, fix $\varepsilon>0$. By continuity of $T$, there exists $\eta>0$ such that for any $H\in\ell^{\infty}(\mathcal{F})$, $\|G-H\|_{\mathcal{F}}<\eta$ implies $|T(G)-T(H)|<\varepsilon$. Let $\rho_{P}(f)=P[(f-Pf)^{2}]^{1/2}$. Since $\mathcal{F}$ is totally bounded (see Van der Vaart and Wellner, 2023, pp. 138-139), there exist $K\geq 1$ and $(f_{1},...,f_{K})\in\mathcal{F}^{K}$ such that $\forall f\in\mathcal{F}$, $\min_{k=1,...,K}\rho_{P}(f,f_{k})<\eta$. Let $\widetilde{G}(f)=G(\widetilde{f})$, where $\widetilde{f}=\arg\min_{g\in\{f_{1},...,f_{K}\}}\rho_{P}(f-g)$. Then,
$|(G-\widetilde{G})f|=|P(f-\widetilde{f}-P(f-\widetilde{f}))|\leq\rho_{P}(f-\widetilde{f})<\eta.$
Hence, $\|G-\widetilde{G}\|_{\mathcal{F}}<\eta$, which implies $|T(G)-T(\widetilde{G})|<\varepsilon$. As a result,
$P(G\psi\in C|\;T(G)\leq 0)\geq\frac{P(G\psi\in C,\;T(\widetilde{G})\leq-\varepsilon)}{H(0)}.$
We can write $T(\widetilde{G})$ as $\widetilde{T}_{K}(Gf_{1},...,Gf_{K})$ for some convex function $\widetilde{T}_{K}$. Let $\mu$ denote the Gaussian distribution of $(G\psi,Gf_{1},...,Gf_{K})$. $\mu$ is a centered Gaussian distribution. Let $E=C\times\mathbb{R}^{K}$ and $F=\mathbb{R}^{p}\times\widetilde{T}^{-1}((-\infty,q-\varepsilon])$. Both $E$ and $F$ are convex. Moreover, they are symmetric about the origin since $T(P)=T(-P)$. Then, by the Gaussian correlation inequality (Royen, 2014),
$P(G\psi\in C,\;T(\widetilde{G})\leq-\varepsilon)=$ $\,\mu(E\cap F)$
$\geq$ $\,\mu(E)\mu(F)$
$=$ $\,P(Z\in C)P(T(\widetilde{G})\leq-\varepsilon),$
where the last equality follows by definition of $Z$. Hence,
$P(G\psi\in C|\;T(G)\leq 0)\geq$ $P(Z\in C)\frac{P(T(\widetilde{G})\leq-\varepsilon)}{H(0)}$
$\geq$ $P(Z\in C)\frac{P(T(G)\leq-2\varepsilon)}{H(0)}$
$\geq$ $P(Z\in C)\frac{H(-2\varepsilon)}{H(0)}.$
Since $H$ is continuous at 00 and $\varepsilon$ was arbitrary, we finally obtain
$P(G\psi\in C|\;T(G)\leq 0)\geq P(Z\in C).$
The result follows by combining the last display with (5).
A.2 Proof of Corollary 1
Let =$\mathcal{E}=\left\{x:\,x^{\prime}V^{-1}x\leq q_{1-\alpha}(p)\right\}$. By assumption, under the null hypothesis $b_{0}=\beta_{0}$,
$F_{n}(b_{0})=n(\widehat{\beta}-\beta_{0})^{\prime}V^{-1}(\widehat{\beta}-\beta_{0})+o_{P}(1).$
As a result,
$P\left(F_{n}(b_{0})\leq q_{1-\alpha}(p)|T_{1,n}\leq q_{1,n},...,T_{J,n}\leq q_{J,n}\right)-P\left(\sqrt{n}(\widehat{\beta}-\beta_{0})\in\mathcal{E}|T_{1,n}\leq q_{1,n},...,T_{J,n}\leq q_{J,n}\right)\to 0$
The result follows by Theorem 3.
A.3 Proof of Theorem 2
The proof is almost the same as that of Theorem 3, and we just highlight the differences. First, since $T_{j}$ is a convex function on $\mathbb{R}^{q}$, it is continuous. Hence, $T(x):=\max_{j=1,...,J}[T_{j}(x)-q_{j}]$ is continuous (and convex). Then, reasoning as above, we obtain
$\lim_{n\to\infty}P\left[\widehat{\Sigma}_{\beta}^{-1/2}\left(\widehat{\beta}-\beta_{0}\right)\in C|T_{n}\leq 0\right]=P(Z\in C|T(Z^{\prime})\leq 0),$
where $T_{n}:=\max_{j=1,...,J}[T_{j,n}-q_{j,n}]$ and
$(Z,Z^{\prime})\sim\mathcal{N}\left(0,\begin{pmatrix}I_{p}&\Sigma_{12}\\ \Sigma_{12}&I_{q}\end{pmatrix}\right).$
The result follows by the Gaussian correlation inequality.